OpenStudy (anonymous):

10/16x-72 x 16x^3-72x^2/8x^2 @amistre64

OpenStudy (anonymous):

\[\frac{ 10 }{ 16x-72 } \times \frac{ 16x ^{3}-72x ^{2} }{8x ^{2} }\]